How much work is required to stop an electron (m=9.11x10^31kg) which is moving with a speed of 1.30x10^6m/s ?
Answer #1

work done W = change of eenrgy

W = 0.5 *mv^2 - 0

W = 0.5*9.11*10^-31 * 1.3*1.3*10^12

W = 7.69 *10^-19 Joules
Answer #2

work done W = change of eenrgy

W = 0.5 *mv^2 - 0

W = 0.5*9.11*10^-31 * 1.3*1.3*10^12

W = 7.69 *10^-19 Joules or 4.806 eV